# AGAV-Rater: Adapting Large Multimodal Model for AI-Generated Audio-Visual Quality Assessment

### v1

@@ Abstract
Many video-to-audio (VTA) methods have been proposed for dubbing silent AI-generated videos. An efficient quality assessment method for AI-generated audio-visual content (AGAV) is crucial for ensuring audio-visual quality. Existing audio-visual quality assessment methods struggle with unique distortions in AGAVs, such as unrealistic and inconsistent elements. To address this, we introduce AGAVQA, the first large-scale AGAV quality assessment dataset, comprising $3,382$ AGAVs from $16$ VTA methods. AGAVQA includes two subsets: AGAVQA-MOS, which provides multi-dimensional scores for audio quality, content consistency, and overall quality, and AGAVQA-Pair, designed for optimal AGAV pair selection. We further propose AGAV-Rater, a LMM-based model that can score AGAVs, as well as audio and music generated from text, across multiple dimensions, and selects the best AGAV generated by VTA methods to present to the user. AGAV-Rater achieves state-of-the-art performance on AGAVQA, Text-to-Audio, and Text-to-Music datasets. Subjective tests also confirm that AGAV-Rater enhances VTA performance and user experience. The project page is available at https://agav-rater.github.io.
@@ 1 Introduction
Traditional audio-visual quality assessment (AVQA) methods (Cao et al., 2023a, b; Min et al., 2020) focus on distortions caused during the capture and transmission stages, which make it difficult to identify unique distortions in AIGC, such as inconsistent A/V content, and unnatural audio. With the rise of LMMs, researchers (Wang et al., 2024b, c; Wu et al., ) utilize the powerful content and language comprehension capabilities of LMMs to evaluate the quality of AIGC images and videos more accurately. However, most quality assessment research focuses on the visual capabilities of LMMs, with little exploration of their A/V capabilities. This raises the question:
@@
In this paper, to develop and refine methods for evaluating AGAV quality, we construct the first AI-generated audio-visual quality assessment dataset, AGAVQA, which contains $3,382$ AGAVs generated by $16$ VTA methods. Fig. 2 illustrates the dataset construction pipeline. We hope that AGAV quality assessment methods can score AGAVs from multiple dimensions, while also assisting VTA models in selecting the optimal result from multiple generated outputs. To meet the above two needs, our AGAVQA dataset is divided into two subsets. In the AGAVQA-MOS subset, we utilize $8$ VTA methods to generate $3,088$ AGAV from $386$ AIGC videos. Then we conduct a subjective experiment to collect $9,264$ Mean Opinion Scores (MOSs) across three dimensions, including audio perceptual quality, A/V content consistency, and overall A/V quality. In the AGAVQA-Pair subset, we collect $294$ AGAVs from $8$ VTA Github pages. The same video content forms a group, resulting in $75$ question-answer pairs. We evaluate $7$ closed-source LMMs on the AGAVQA-Pair subset to assess accuracy in selecting the optimal AGAVs, as shown in Fig. 1. The results indicate that LMMs still have significant room for improvement in evaluating AGAV quality. Moreover, existing closed-source LMMs have difficulty providing a numerical score for AGAV quality as humans do. Therefore, this paper primarily focuses on:
@@
Our experimental results demonstrate that AGAV-Rater achieves state-of-the-art performance on three quality assessment datasets: AGAVQA-MOS, text-to-audio (TTA), and text-to-music (TTM) (Deshmukh et al., 2024). Since the video content and VTA methods in the AGAVQA-MOS and AGAVQA-Pair subsets do not overlap, we validate the generalization ability and robustness of AGAV-Rater on the unseen dataset AGAVQA-Pair, as shown in Fig. 1. We further conduct a subjective experiment and find that AGAV-Rater helps VTA methods select high-quality generated results to present to users, thereby enhancing user experience. Our core contributions can be summarized as three-fold:
  • A large-scale AGAV quality assessment dataset AGAVQA. It label AGAVs quality in two ways: multi-dimensional numerical scores and the optimal AGAV pair.
@@
Figure 2: The construction process of the AGAVQA dataset. The AGAVQA dataset is divided into two subsets: AGAVQA-MOS and AGAVQA-Pair, which involve multi-dimensional score prediction and optimal AGAV pair selection tasks.
@@ 2.1 Audio-Visual Quality Assessment Dataset
Early research focused on compression distortions during transmission, leading to the construction of several traditionally distorted AVQA datasets. The largest one is the LIVE-SJTU dataset proposed by Min et al. (Min et al., 2020), which contains $14$ original high-quality A/V consequences and $336$ degraded ones. With the development of streaming media, researchers (Cao et al., 2023b; Ying et al., 2022) found that user-uploaded videos have more complex and diverse distortions, thus constructing authentically distorted AVQA datasets. With the rise of generative models, AIGC exhibits unique distortions that do not occur in real-world scenarios. To explore the distortions and perceived quality of AGAVs, we establish the first AGAV quality assessment dataset, AGAVQA. Compared to single-modal AIGC image (video) quality assessment datasets, AGAVQA tackles more complex multimodal challenges, such as A/V content inconsistency and synchronization issues.
@@ 2.2 Quality Assessment Methods
Audio-Visual Quality Assessment. Compared to AQA, AVQA is a more complex task as it requires handling the interaction between video and audio modalities. Early research (Min et al., 2020) utilized Support Vector Regression (SVR) to predict A/V quality scores by regressing handcrafted features extracted from video and audio. Although this method is effective for evaluating compressed A/V content, it performs poorly in real-world scenarios with mixed distortions. To address this issue, researchers (Cao et al., 2023b) have proposed deep learning-based approaches to predict A/V quality in real-world environments. However, these methods are not suitable for AGAVs, as they do not consider A/V content consistency and fail to address unnatural audios often present in AIGC scenarios. Our proposed AGAV-Rater not only predicts the quality of AGAVs but also evaluates the quality of TTAs and TTMs. We validate the effectiveness of AGAV-Rater on the AGAVQA, TTA, and TTM datasets.
@@
3 AGAVQA Dataset Construction
Our proposed AGAVQA dataset consists of two subsets: the AGAVQA-MOS and AGAVQA-Pair, designed for multi-dimensional score prediction and optimal AGAV selection, respectively. In this section, we introduce the construction process of two subsets and analyze the subjective scores in the AGAVQA-MOS subset, as illustrated in Fig. 2.
@@ 3.1 AGAVQA-MOS Subset Generation
AIGC Video Collection. We first collected high-quality AIGC videos from AIGC video display websites (Sora (sor, 2024), KLing (kli, 2024), and Gen3 (Gen, 2024)) and the video generation benchmark Vbench (Huang et al., 2024). Additionally, we gathered prompts from FETV (Liu et al., 2024b) and generated AIGC videos using closed-source text-to-video platforms (Pika 1.0 (pik, 2023) and Gen3 (Gen, 2024)). From these AIGC videos, we manually selected $386$ high-quality videos with clear audio source information, covering $11$ types of audio sources. The distribution of main audio sources in AIGC videos is shown in step $1$ of the AGAVQA-MOS construction process in Fig. 2.
@@ 3.2 MOSs Analysis
In the column labeled “Comparison of MOS Across $8$ VTA Methods in AGAVQA-MOS” in Fig. 2, the maximum, minimum, and average subjective scores for the $8$ VTA methods are presented across the three dimensions. We can observe that the AGAVs generated by SVA (Chen et al., 2024) exhibit the best audio quality, which can be attributed to the use of proprietary TTA tools, i.e., AudioGen (Kreuk et al., 2022) and MusicGen (Copet et al., 2024), to generate high-quality sound effects and background music. ElevenLabs achieves the best A/V content consistency and overall quality by using ChatGPT-4 to extract key sound information from videos and generate high-quality audio with its proprietary TTA tool, ensuring seamless coherence between audio and video.
@@ 3.3 AGAVQA-Pair Subset Collection
Most VTA Github pages exhibit superior performance by dubbing the same video using both their approach and other VTA methods. We collected $294$ publicly available AGAVs from $8$ VTA Github pages, including SSV2A (Guo et al., 2024a), ReWaS (Jeong et al., 2025), TIVA (Wang et al., 2024d), V2A-Mapper (Wang et al., 2024a), STAV2A (Ren et al., 2024), V2A-SceneDetector (Yi & Li, 2024), Frieren (Wang et al., 2024g), and SonicVisionLM (Xie et al., 2024). AGAVs with the same video content are grouped, with the optimal AGAVs already labeled on the Github pages. We manually verified the accuracy of the optimal AGAVs, and then formed $75$ instruction-response pairs as follows:
@@
Due to the inherent subjectivity in human evaluations of AGAVs, different subjects may provide varying scores to the same AGAV. However, when selecting the optimal AGAV, most subjects tend to give the same choice, leading to more reliable quality labels. In some application scenarios, businesses only need to select the highest-quality result from multiple generated options to present to users, without requiring detailed quality scores for each AGAV. Therefore, we construct the AGAVQA-Pair subset. This subset evaluates the performance of AGAV quality assessment methods by measuring their accuracy in selecting the optimal AGAV. More details of the AGAVQA dataset are provided in Appendix A.
@@ 4.1 Pre-Training via Text-Defined Levels
To alleviate the burden of constructing large-scale quality assessment datasets, we first construct $50,952$ instruction-response pairs from $3$ real-world audio-caption related datasets for AGAV-Rater pre-training, as shown in Fig. 3. The responses are automatically labeled with two text-defined levels (excellent and bad). We select the VGGSound, AudioCaps, and MusicCaps datasets to cover $3$ different scenarios: audio-video, audio-text, and music-text. Instruction-response pairs are designed from two perspectives: content consistency and audio quality. Take the A/V scenario as an example, the instruction for content consistency related pairs is as follows:
@@ 4.3 Model Structure
Question Type: Multi-Input Comparison
@@ 5.1 Experimental Settings
In this paper, we fine-tune the AGAV-Rater from the pre-trained weights of VideoLLaMA2 (Cheng et al., 2024). The AGAV-Rater model is implemented with PyTorch and trained on two 96GB H20 GPUs. The learning rate is set to $51e-$ 5, and the batch size is set to $9$. During pre-training, the number of training epochs is set to $1$. For fine-tuning, the number of training epochs is set to $5$ on the AGAVQA-MOS subset and $10$ on the TTA and TTM datasets. All experiments for each method are retrained on the AGAVQA-MOS subset using 5-fold cross-validation. The reported performance of the AGAV-Rater is evaluated on the final weights after training.
@@ 5.2 Compared Methods
Except for audio-visual LMMs, all methods are retrained on the AGAVQA-MOS, TTA, and TTM datasets after loading the default weights. Audio-visual LMMs are directly tested on the dataset using their default weights. For quality-related questions, most audio-visual LMMs respond with text-defined quality levels and are unable to stably output numerical scores. Therefore, following the testing scheme designed in (Wu et al., ), we prompt LMMs to receive quality level tokens. Then, extract the probability distribution $\mathcal{X}$ of the predicted level token and convert it into the final predicted scores $\mathrm{S_{LMM}}$ as follows:
@@ 5.4 Evaluation on Optimal AGAV Selection Tasks
Since there is no overlap between the video content in the AGAVQA-MOS and AGAVQA-Pair subsets, we perform cross-dataset validation experiments by training the AGAV-Rater on the AGAVQA-MOS subset and testing on the unseen AGAVQA-Pair subset. We collect $11$ audio-visual LMMs as comparison methods, including $4$ open-source LMMs and $7$ closed-source LMMs. For GPT-4o$+$audio caption, we utilize GPT-4o-audio to generate captions for audio, then feed the text, video, and audio captions into GPT-4o. Similarly, for GPT-4o$+$video caption, we utilize GPT-4o to generate captions for video, then feed the text, audio, and video captions into GPT-4o-audio. As shown in Tab. 4.3, we test their accuracy in answering optimal AGAV questions on the AGAVQA-Pair subset. We designed two types of instructions to prompt LMMs for the optimal AGAV. For multi-input comparisons instruction type, we utilize instruction-response pairs in the AGAVQA-Pair subset to let LMMs answer the optimal audio number. To prevent the model from guessing, we shuffle the audio order and iterate through each audio number as the correct answer, inputting them into LMMs to calculate the response accuracy. This instruction type applies to both open-source and closed-source LMMs, with optimal AGAV judgment based on their textual responses. For the single-input scoring instruction type, we sequentially ask the overall quality of each AGAV and compute the final predicted scores by Eq. 1. The AGAV with the highest score is selected as the optimal AGAV. Since closed-source LMMs cannot access the token probability distribution, this method is only applied to open-source LMMs. As seen in Tab. 4.3, our method achieves the highest accuracy across all $8$ VTA methods, demonstrating that our model has the best ability to judge the optimal AGAV and exhibits strong transferability.

### v2

@@ Abstract
Many video-to-audio (VTA) methods have been proposed for dubbing silent AI-generated videos. An efficient quality assessment method for AI-generated audio-visual content (AGAV) is crucial for ensuring audio-visual quality. Existing audio-visual quality assessment methods struggle with unique distortions in AGAVs, such as unrealistic and inconsistent elements. To address this, we introduce AGAVQA-3k, the first large-scale AGAV quality assessment dataset, comprising $3,382$ AGAVs from $16$ VTA methods. AGAVQA-3k includes two subsets: AGAVQA-MOS, which provides multi-dimensional scores for audio quality, content consistency, and overall quality, and AGAVQA-Pair, designed for optimal AGAV pair selection. We further propose AGAV-Rater, a LMM-based model that can score AGAVs, as well as audio and music generated from text, across multiple dimensions, and selects the best AGAV generated by VTA methods to present to the user. AGAV-Rater achieves state-of-the-art performance on AGAVQA-3k, Text-to-Audio, and Text-to-Music datasets. Subjective tests also confirm that AGAV-Rater enhances VTA performance and user experience. The dataset and code is available at https://github.com/charlotte9524/AGAV-Rater.
@@ 1 Introduction
Traditional audio-visual quality assessment (AVQA) methods (Cao et al., 2023a, b; Min et al., 2020) focus on distortions caused during the capture and transmission stages, which make it difficult to identify unique distortions in AIGC, such as inconsistent A/V content, and unnatural audio. With the rise of LMMs, researchers (Wang et al., 2024b, c; Wu et al., ) utilize the powerful content and language comprehension capabilities of LMMs to evaluate the quality of AIGC images and videos more accurately. However, most quality assessment research focuses on the visual capabilities of LMMs (Sun et al., 2024b, a, 2023), with little exploration of their A/V capabilities. This raises the question:
@@
In this paper, to develop and refine methods for evaluating AGAV quality, we construct the first AI-generated audio-visual quality assessment dataset, AGAVQA-3k, which contains $3,382$ AGAVs generated by $16$ VTA methods. Fig. 2 illustrates the dataset construction pipeline. We hope that AGAV quality assessment methods can score AGAVs from multiple dimensions, while also assisting VTA models in selecting the optimal result from multiple generated outputs. To meet the above two needs, our AGAVQA-3k dataset is divided into two subsets. In the AGAVQA-MOS subset, we utilize $8$ VTA methods to generate $3,088$ AGAV from $386$ AIGC videos. Then we conduct a subjective experiment to collect $9,264$ Mean Opinion Scores (MOSs) across three dimensions, including audio perceptual quality, A/V content consistency, and overall A/V quality. In the AGAVQA-Pair subset, we collect $294$ AGAVs from $8$ VTA Github pages. The same video content forms a group, resulting in $75$ question-answer pairs. We evaluate $7$ closed-source LMMs on the AGAVQA-Pair subset to assess accuracy in selecting the optimal AGAVs, as shown in Fig. 1. The results indicate that LMMs still have significant room for improvement in evaluating AGAV quality. Moreover, existing closed-source LMMs have difficulty providing a numerical score for AGAV quality as humans do. Therefore, this paper primarily focuses on:
@@
Our experimental results demonstrate that AGAV-Rater achieves state-of-the-art performance on three quality assessment datasets: AGAVQA-MOS, text-to-audio (TTA), and text-to-music (TTM) (Deshmukh et al., 2024). Since the video content and VTA methods in the AGAVQA-MOS and AGAVQA-Pair subsets do not overlap, we validate the generalization ability and robustness of AGAV-Rater on the unseen dataset AGAVQA-Pair, as shown in Fig. 1. We further conduct a subjective experiment and find that AGAV-Rater helps VTA methods select high-quality generated results to present to users, thereby enhancing user experience. Our core contributions are threefold:
  • A large-scale AGAV quality assessment dataset AGAVQA-3k. It labels AGAVs’ quality in two ways: multi-dimensional numerical scores and the optimal AGAV pair.
@@
Figure 2: The construction process of the AGAVQA-3k dataset. The AGAVQA-3k dataset is divided into two subsets: (a) AGAVQA-MOS and (b) AGAVQA-Pair, which involve multi-dimensional score prediction and optimal AGAV pair selection tasks. (c) We present the maximum, minimum, and average subjective scores for the $8$ VTA methods across the three dimensions.
@@ 2.1 Audio-Visual Quality Assessment Dataset
Early research focused on compression distortions during transmission, leading to the construction of several traditionally distorted AVQA datasets. The largest one is the LIVE-SJTU dataset proposed by Min et al. (Min et al., 2020), which contains $14$ original high-quality A/V consequences and $336$ degraded ones. With the development of streaming media, researchers (Cao et al., 2023b; Ying et al., 2022) found that user-uploaded videos have more complex and diverse distortions, thus constructing authentically distorted AVQA datasets. With the rise of generative models, AIGC exhibits unique distortions that do not occur in real-world scenarios. To explore the distortions and perceived quality of AGAVs, we establish the first AGAV quality assessment dataset, AGAVQA-3k. Compared to single-modal AIGC image (video) quality assessment datasets, AGAVQA-3k dataset tackles more complex multimodal challenges, such as A/V content inconsistency and synchronization issues.
@@ 2.2 Quality Assessment Methods
Audio-Visual Quality Assessment. Compared to AQA, AVQA is a more complex task as it requires handling the interaction between video and audio modalities. Early research (Min et al., 2020) utilized Support Vector Regression (SVR) to predict A/V quality scores by regressing handcrafted features extracted from video and audio. Although this method is effective for evaluating compressed A/V content, it performs poorly in real-world scenarios with mixed distortions. To address this issue, researchers (Cao et al., 2023b) have proposed deep learning-based approaches to predict A/V quality in real-world environments. However, these methods are not suitable for AGAVs, as they do not consider A/V content consistency and fail to address unnatural audios often present in AIGC scenarios. Our proposed AGAV-Rater not only predicts the quality of AGAVs but also evaluates the quality of TTAs and TTMs. We validate the effectiveness of AGAV-Rater on the AGAVQA-3k, TTA, and TTM datasets.
@@
3 Dataset Construction
Our proposed AGAVQA-3k dataset consists of two subsets: the AGAVQA-MOS and AGAVQA-Pair, designed for multi-dimensional score prediction and optimal AGAV selection, respectively. In this section, we introduce the construction process of two subsets and analyze the subjective scores in the AGAVQA-MOS subset, as illustrated in Fig. 2.
@@ 3.1 AGAVQA-MOS Subset Generation
AIGC Video Collection. We first collected high-quality AIGC videos from AIGC video display websites (Sora (sor, 2024), KLing (kli, 2024), and Gen3 (Gen, 2024)) and the video generation benchmark Vbench (Huang et al., 2024). Additionally, we gathered prompts from FETV (Liu et al., 2024b) and generated AIGC videos using closed-source text-to-video platforms (Pika 1.0 (pik, 2023) and Gen3 (Gen, 2024)). From these AIGC videos, we manually selected $386$ high-quality videos with clear audio source information, covering $11$ types of audio sources. The distribution of main audio sources in AIGC videos is shown in Fig. 2(a).
@@ 3.2 MOSs Analysis
In Fig. 2(c), the maximum, minimum, and average subjective scores for the $8$ VTA methods are presented across the three dimensions. We can observe that the AGAVs generated by SVA (Chen et al., 2024) exhibit the best audio quality, which can be attributed to the use of proprietary TTA tools, i.e., AudioGen (Kreuk et al., 2022) and MusicGen (Copet et al., 2024), to generate high-quality sound effects and background music. ElevenLabs achieves the best A/V content consistency and overall quality by using ChatGPT-4 to extract key sound information from videos and generate high-quality audio with its proprietary TTA tool, ensuring seamless coherence between audio and video. Krippendorff’s $\alpha$ (Hayes & Krippendorff, 2007) can be used to measure the quality of the subjects’ ratings. We calculate Krippendorff’s $\alpha$ for audio quality, A/V content consistency, and A/V overall quality, which are $0.6814$, $0.7343$, and $0.7143$, respectively, indicating appropriate variations among subjects. We also randomly divide subjects into two groups and calculate the SRCC of average scores between the two groups. After ten repetitions, the average SRCC for audio quality, A/V content consistency, and A/V overall quality are $0.8043$, $0.8318$, and $0.8297$, validating rating consistency.
@@ 3.3 AGAVQA-Pair Subset Collection
Most VTA Github pages exhibit superior performance by dubbing the same video using both their approach and other VTA methods. As shown in Fig. 2(b), we collected $294$ publicly available AGAVs from $8$ VTA Github pages, including SSV2A (Guo et al., 2024a), ReWaS (Jeong et al., 2025), TIVA (Wang et al., 2024d), V2A-Mapper (Wang et al., 2024a), STAV2A (Ren et al., 2024), V2A-SceneDetector (Yi & Li, 2024), Frieren (Wang et al., 2024g), and SonicVisionLM (Xie et al., 2024). These AGAVs, sourced from third-party platforms, offer a more objective and impartial dataset. These VTA Github pages are all released in the past year, representing the latest technology in VTA methods. AGAVs with the same video content are grouped, with the optimal AGAVs already labeled on the Github pages. We manually verified the accuracy of the optimal AGAVs, and then formed $75$ instruction-response pairs as follows:
@@
Due to the inherent subjectivity in human evaluations of AGAVs, different subjects may provide varying scores to the same AGAV. However, when selecting the optimal AGAV, most subjects tend to give the same choice, leading to more reliable quality labels. In some application scenarios, businesses only need to select the highest-quality result from multiple generated options to present to users, without requiring detailed quality scores for each AGAV. Therefore, we construct the AGAVQA-Pair subset. This subset evaluates the performance of AGAV quality assessment methods by measuring their accuracy in selecting the optimal AGAV. More details of the AGAVQA-3k dataset are provided in Appendix A.
@@ 4.1 Pre-Training via Text-Defined Levels
To alleviate the burden of constructing large-scale quality assessment datasets, we first construct $50,952$ instruction-response pairs from $3$ real-world audio-caption related datasets for AGAV-Rater pre-training, as shown in Fig. 3. The responses are automatically labeled with two text-defined levels (excellent and bad). We select the VGGSound, AudioCaps, and MusicCaps datasets to cover $3$ different scenarios: audio-video, audio-text, and music-text. In the $50,952$ instruction-response pairs, the audio-video, audio-text, and music-text scenarios contain $25,592$, $19,000$, and $6,000$ pairs, respectively. Instruction-response pairs are designed from two perspectives: content consistency and audio quality. Under each scenario, half of the pairs focus on content consistency, and the other half on audio quality. Take the A/V scenario as an example, the instruction for content consistency related pairs is as follows:
@@ 4.3 Model Structure
Random 0.33 0.20 0.41 0.20 0.50 0.25 0.20 0.25 0.32
[…]
@@ 5.1 Experimental Settings
In this paper, we fine-tune the AGAV-Rater from the pre-trained weights of VideoLLaMA2 (Cheng et al., 2024). The AGAV-Rater model is implemented with PyTorch and trained on two 96GB H20 GPUs. The learning rate is set to $51e-$ 5, and the batch size is set to $9$. During pre-training, the number of training epochs is set to $1$, and optimization is performed. For fine-tuning, the number of training epochs is set to $5$ on the AGAVQA-MOS subset and $10$ on the TTA and TTM datasets (Deshmukh et al., 2024). Fine-tuning the AGAV-Rater model on the AGAVQA-MOS subset for $5$ epochs using two 96GB H20 GPUs takes approximately $5$ hours. All experiments for each method are retrained on the AGAVQA-MOS subset using 5-fold cross-validation. The reported performance of the AGAV-Rater is evaluated on the final weights after training.
@@ 5.2 Compared Methods
Except for audio-visual LMMs, all methods are retrained on the AGAVQA-MOS, TTA, and TTM datasets after loading the default weights. Original multimodal alignment methods extract audio and video features using their encoders, then align them into a common vector space. We load the default parameters and use these encoders to extract audio and video features, which are then concatenated. The concatenated features are subsequently fed into a fully connected layer with an output dimension of $3$ to predict the three-dimensional scores.
Audio-visual LMMs are directly tested on the dataset using their default weights. For quality-related questions, most audio-visual LMMs respond with text-defined quality levels and are unable to stably output numerical scores. Therefore, following the testing scheme designed in (Wu et al., ), we prompt LMMs to receive quality level tokens. Then, extract the probability distribution $\mathcal{X}$ of the predicted level token and convert it into the final predicted scores $\mathrm{S_{LMM}}$ as follows:
@@
Table 5: Ablation study of the base models.
Dimension Audio Quality Consistency Overall Quality
Metric SRCC PLCC SRCC PLCC SRCC PLCC
GroundingGPT (ACL 2024) 0.4387 0.4494 0.5067 0.4764 0.4975 0.5297
OneLLM (CVPR 2024) 0.6578 0.6879 0.6184 0.6297 0.6327 0.6388
AGAV-Rater 0.7909 0.8108 0.7553 0.7645 0.7458 0.7552
@@ 5.4 Evaluation on Optimal AGAV Selection Tasks
The AGAVQA-Pair subset was collected from $8$ VTA webpages, dividing it into $8$ corresponding categories. For each category, the optimal AGAV is generated by the corresponding VTA method. Since there is no overlap between the video content in the AGAVQA-MOS and AGAVQA-Pair subsets, we perform cross-dataset validation experiments by training the AGAV-Rater on the AGAVQA-MOS subset and testing on the unseen AGAVQA-Pair subset. We collect $11$ audio-visual LMMs as comparison methods, including $4$ open-source LMMs and $7$ closed-source LMMs. For GPT-4o$+$audio caption, we utilize GPT-4o-audio to generate captions for audio, then feed the text, video, and audio captions into GPT-4o. Similarly, for GPT-4o$+$video caption, we utilize GPT-4o to generate captions for video, then feed the text, audio, and video captions into GPT-4o-audio. As shown in Tab. 4.3, we test their accuracy in answering optimal AGAV questions on the AGAVQA-Pair subset. We designed two types of instructions to prompt LMMs for the optimal AGAV. For multi-input comparisons instruction type, we utilize instruction-response pairs in the AGAVQA-Pair subset to let LMMs answer the optimal audio number. To prevent the model from guessing, we shuffle the audio order and iterate through each audio number as the correct answer, inputting them into LMMs to calculate the response accuracy. This instruction type applies to both open-source and closed-source LMMs, with optimal AGAV judgment based on their textual responses. For the single-input scoring instruction type, we sequentially ask the overall quality of each AGAV and compute the final predicted scores by Eq. 1. The AGAV with the highest score is selected as the optimal AGAV. Since closed-source LMMs cannot access the token probability distribution, this method is only applied to open-source LMMs. The above $11$ audio-visual LMMs use their original model parameters without fine-tuning on the AGAVQA-MOS subset. We also test the accuracy of the audio-video alignment methods which have been fine-tuned on the AGAVQA-MOS subset. As seen in Tab. 4.3, our method achieves the highest accuracy across all $8$ VTA methods, demonstrating that our model has the best ability to judge the optimal AGAV and exhibits strong transferability.